# Physics of Strong Magnetism with eXTP

### v1

@@ 2.2.2 Magnetized atmosphere and magnetic condensation
There is no doubt that the surfaces of accreting NSs in close binary systems are covered with gaseous envelopes. As a result, the emergent radiation originates from the upper layers of this envelope, which can be called the atmosphere [see, for example, the atmospheres of X-ray bursters 118]. However, the question of the presence of gaseous atmospheres also around isolated NSs remained open for a long time. Observational data support the idea that emergent radiation may be formed in an atmosphere of at least one class of isolated NSs, the central compact objects (CCOs), young sources still located near the center of their supernova remnants [149]. Probably, the envelopes of these NSs have formed as a result of fallback accretion just after their birth [36].
@@
The atmospheres of more highly magnetized NSs ( G are more challenging to model, but have been studied in connection with magnetars to explain the thermal blackbody-like component detected in their emission spectra [see e.g. 75, 143, 240, 250, see also Potekhin 157 for a review].
@@
The main effect is suppression of the X-mode opacity by a factor [218]. Therefore, a typical feature of magnetic atmospheres for which is the presence of two photospheres, one for X-mode and the other for O-mode photons. The X-mode photosphere is located deeper in the atmosphere, and therefore, at least if the NSs is modeled as a passive cooler without heat deposition in its external layers, its temperature is higher than that of the O-mode photosphere. As a result, such an atmosphere should radiate mainly in the X-mode, and its emergent radiation should be almost completely linearly polarized.
@@
However, the possibility that the radiating surface of magnetars is described by a classical semi-infinite atmosphere without its own energy sources is still strongly debated [see e.g. 197]. Possibly, energy can be released throughout the entire thickness of the atmosphere as a result of magnetic field dissipation. Heat deposition in the outer layers by returning magnetospheric currents can also largely affect polarization [64, 97]. eXTP has the revolutionary capability to observe the polarization of thermal radiation from a large number of magnetars, extending the preliminary finding by IXPE, and this will unveil the properties of the radiating surface responsible for the thermal component in the spectra of magnetars.
In the presence of a strong magnetic field, the NS atmosphere may turn into a condensed state. This occurs for sufficiently low temperatures, at which atoms, elongated along the direction of , form molecular chains via covalent bonding [110, 108, 215, 159]. Such elongated atoms settle on the star’s surface in a solid/liquid phase, leaving the NS surface exposed (a phenomenon called “magnetic condensation”). There is in fact theoretical and observational evidence that this may occur in some of the colder and more magnetized NSs, such as magnetars (see the next section and references therein) and possibly the XDINSs (the so-called “Magnificent Seven”) [78, 20]. The discrimination between atmospheric and condensed surface emission can be easily performed by eXTP, since low polarization degree is expected for radiation coming from a condensed surface [, dominated by either O- or X-mode photons according to the geometry of view, 63, 197]
@@
Another exciting potentiality of eXTP would be the detection of polarization of X-ray emission from CCOs, since it could resolve the problem of the chemical composition of the atmospheres of some of these non-pulsating objects, such as those in the supernova remnants Cas A [146] and HESS J1731–347 [103]. The assumption of temperature uniformity on the surface of these NSs led to the conclusion that their atmospheres are made of carbon [74], and, in the case of HESS J1731–347, led to the estimate of an extremely low mass of the NS, [44]. An alternative and intriguing scenario is that these NSs host instead very high magnetar-like magnetic fields buried by fall-back accretion [see e.g. 19], in which case their surfaces may exhibit hot spots covered by hydrogen atmospheres. In that case, the absence of pulsations may be attributed to a very unfavorable viewing geometry [185]. As discussed in [187], for carbon atmospheres the expected polarization is not detectable, while detecting even a modest degree of polarization (up to a few percent for the CCO in HESS J1731-347, in the 2-8 keV band) would support the hot spot hydrogen atmosphere hypothesis. These authors estimated that measuring this polarization with IXPE required a long exposure time of a few million of seconds, while eXTP, with his enhanced sensitivity, would be excellently placed to make the detection.
@@ 2.2.3 Spectral-continuum polarimetry in magnetars
The nature of the thermal component remains an open problem. Given the ultra-strong magnetic fields of these objects, emission from the star surface may either emerge from a magnetized atmosphere (possibly bombarded by the returning magnetospheric currents) or from a “bare” condensed surface. Despite the fact that the spectrum emitted in the two cases is expected to be not much different (and close to a blackbody distribution), polarization properties are poles apart. As mentioned above, a much lower polarization degree is characteristic of radiation from a condensed surface, compared to the case of a magnetized atmosphere, for which a high polarization degree is generally expected, even when QED mode conversion is accounted for. Because for magnetar-like magnetic fields (), the energy at which mode switching occurs is outside the – band, the working energy range for GPD-based polarimeters like IXPE and eXTP [96].
@@
Different behaviors have been observed in the AXPs 1RXS J170849.0400910 and 1E 1841045. Although the relatively low degree of polarization at low energies (–) still argues in favor of a condensed surface origin, it increases monotonically with energy, up to – at –. Moreover, contrary to the case of 4U 014261, the polarization angle is basically constant throughout the – range. In 1E 1841045, a simultaneous observation with NuSTAR allowed us to associate the high degree of polarization at high energies with a power-law component, which still dominates at [169, 181]. Such a broad-band analysis was not possible for 1RXS J170849.0400910, making it difficult to understand whether the observed polarization properties can only be attributed to surface / magneticospheric emission. Once more, eXTP observations of these sources (with both the SFA and PFA instruments) will provide invaluable insight, helping to fully characterize the spectral continuum of magnetar sources.
@@ 2.3.1 RVM and beyond
Using the RVM, one can obtain important information about the source geometry, mainly the magnetic obliquity (i.e. the angle between the magnetic axis and rotational axis) and the inclination angle between the line of sight and rotational axis, as well as the position angle of the rotational axis projected onto the plane of sky. In each phase, the observer sees different position angles (PAs, position angle of the rotating “vector” relative to the meridian plane defined by the rotational axis and line of sight). Therefore, compared to the case of a NS endowed with a dipolar magnetic field, the RVM will be modified if (i) the “vector” intrinsically changes with time (e.g. due to forced or free precession) or (ii) the field topology deviates from a dipole, e.g. because of the presence of a toroidal component (a twist).
The geometry of a rotating dipole field is shown in Fig. 5. Using simple arguments from spherical trigonometry, it is possible to obtain PA as a function of phase [see, e.g., 160]
@@ 2.3.1.1 Change of the geometry due to precession
In practice, a simpler and straightforward approach can be used by fitting the observed polarization angle with the standard RVM model at different epochs. Then, different sets of RVM parameters will be obtained. In the case of forced precession, the rotation axis of the NS experiences a significant change, due to gravity effects produced by a companion star or a fallback disc. Then, the magnetic axis is dragged together with the spin, so that remains constant. In contrast, free precession arises because of an aspherical deformation of the star, possibly produced by the strong -field. In this case the rotation axis is practically fixed in space (which is also coincide with the star’s angular momentum), so that and constant, while the inclination of the magnetic axis varies in time, and a consequence also changes.
@@ 2.3.1.2 Change of field line geometry due to the twist of the magnetic field
If X-ray polarization observations of magnetars are performed during an outburst, a monotonically evolving PA (especially and ) is expected. This is shown in Fig. 7, where the polarization angle resulting from different eXTP simulations is shown [see 196, 205]. Four stages of an outburst have been considered, characterized by different values of the flux – (unabsorbed) and of the radial index of the globally twisted field [see 200]. For simplicity, the source emission is modeled assuming a BB distribution (at a temperature , constant across the surface), polarized in the X-mode. According to the results shown in Fig. 7, the deviation from the classical RVM for different magnetic field twists can be clearly resolved by eXTP observations with sufficiently long exposure.
@@ 2.4.1 Glitches
Starquakes in magnetars might trigger FRBs [222, 219, 190, 225]. Some theoretical and statistical works reveal that FRBs have some Earthquake-like behaviors [206]. The energy released during the starquakes is estimated to support X-ray bursts via resonant inverse Compton scattering [234]. The quake model predicts that FRBs will be associated with glitches and quasi-periodic oscillations (QPOs), which are expected to be found in the associated X-ray bursts [113]. Currently, several FRB-like bursts from SGR J1935+2154 are found to be associated with the source spin changes. [60] found a glitch candidate before FRB 200428, the fast rise of this glitch was followed by a delayed spin-up rise, similar to that observed in the Crab pulsar [176]. It is reported that two neighboring glitches within a time interval of approximately nine hours with great details, and the two glitches bracket an FRB on 2022 October 14 [83]. Rapid spin down phase was observed between the two glitches. Previously, [236] reported a spin-down glitch around 2020 October 5 that occurred before three FRB-like bursts, and the observational cadence seems to be much lower than that in [83]. The possibility that this spin-down glitch is actually a glitch followed by an enhanced spin-down rate (just as what was observed by Hu et al. [83]) is worthy of consideration. Given the apparent association between glitches/spin-down glitches and galactic FRBs, quick responses and high-cadence observations around the epoch of FRBs should be conducted. Research on the physical origin of these glitches/spin-down glitches will help to reveal the physical origin of galactic FRBs eventually.
@@ 2.4.2.1 Bursting activity and outbursts in magnetar sources
One of the most distinctive behaviors of magnetars is the occurrence of events in which huge amounts of energy are released over different timescales, which, in many cases, led to the discovery of new magnetar sources. Persistent magnetars are characterized by the emission of short (–) bursts, with luminosities in the range –, and intermediate (–) flares, with peak luminosities ranging from to [216]. The rarest giant flares, with luminosities up to in the short (–) initial spike and decreasing down to in the pulsating tail (modulated at the spin frequency of the star), have been definitively observed from only three magnetars [SGR 052666, SGR 190014 and SGR 180620, see 128, 85, 86, 145]. Several magnetar giant flare candidates have been reported from nearby galaxies, such as GRB 200415A in NGC 253 [191]. Conversely, outbursts are typically identified in transient sources following the detection of X-ray bursts by the wide-field instruments on-board such as Swift, Fermi, GECAM missions. These events are characterized by a sudden flux enhancement (by a factor of –), followed by a return to quiescence over timescales ranging from months to several years [30]. In particular, the recent launch of the Einstein Probe and SVOM missions has significantly enhanced our ability to detect new magnetar outbursts [237, 226]. With its large FOV, accurate localization and coverage of the hard X-ray band, eXTP/W2C will be powerful to monitor the burst activity of known magnetars and discover new magnetar sources either in our galaxy or nearby galaxies.
@@
Observing an outburst with eXTP, with its capability of performing simultaneous phase-resolved spectral and polarization measurements, will enable the extraction of extensive information regarding aspects such as the activation mechanism of these phenomena, the energy source driving their long-term emission, and the potential correlations with the emission of short bursts. At the onset of the event, a spectral hardening is usually observed, followed by a slow decay phase in which the X-ray spectral shape softens. If outbursts are indeed related to local heat deposition in the outer crust, caused by the release of internal magnetic energy or by the external magnetic field line untwisting, a very peculiar polarization pattern is expected to be associated with the changing spectral shape, allowing us to gain insight into both the magnetic field topology and the physical state of the surface during the event.
In the spectral observations carried out so far, radiation near the peak of an outburst is typically observed to be emitted from large regions, where, due to heat deposition, temperatures are expected to be relatively higher than those of the rest of the surface. This situation represents the most favorable condition to probe the effects of vacuum birefringence (see §2.1). In fact, a sufficiently high surface temperature in the emitting region would prevent magnetic condensation of the gaseous NS atmosphere to occur, so that the intrinsic polarization of emitted radiation can reasonably be expected to be very high. Moreover, if vacuum birefringence were not effective, the observed polarization degree of radiation coming from such an extended and hot cap should be much lower than the intrinsic polarization at the emission, due to the geometrical depolarization caused by tangled magnetic field topology at the surface (see Fig. 2). Hence, measuring a high degree of polarization at the onset of the outburst would provide long-awaited evidence of vacuum birefringence. Furthermore, as the emitting cap shrinks in the decay phase, an increase of the observed polarization degree should be observed if QED effects were not at work, due to geometrical depolarization being less effective for emitting region sizes smaller than (see again Fig. 2). Thanks to the increased effective area and much shorter repointing time compared to those of IXPE, eXTP can map both spectral and polarization properties during a typical magnetar outburst, from the peak phase to the afterglow, so as to put a definitive end to the long-standing problem of the observational test of vacuum birefringence.

### v2

@@ 2.2.2 Magnetized atmosphere and magnetic condensation
There is no doubt that the surfaces of accreting NSs in close binary systems are covered with gaseous envelopes. As a result, the emergent radiation originates from the upper layers of this envelope, which can be called the atmosphere [see, for example, the atmospheres of X-ray bursters 53]. However, the question of the presence of gaseous atmospheres also around isolated NSs remained open for a long time. Observational data support the idea that emergent radiation may be formed in an atmosphere of at least one class of isolated NSs, the central compact objects (CCOs), young sources still located near the center of their supernova remnants [54]. Probably, the envelopes of these NSs have formed as a result of fallback accretion just after their birth [55].
@@
The atmospheres of more highly magnetized NSs ( G are more challenging to model, but have been studied in connection with magnetars to explain the thermal blackbody-like component detected in their emission spectra [see e.g. 65, 66, 67, 68, see also Potekhin 69 for a review].
@@
where , , and are the reduced Plank constant, light speed, charge and mass of an electron while and denote magnetic field and magnetic strength in units of Gauss. The main effect is suppression of the X-mode opacity by a factor [70]. Therefore, a typical feature of magnetic atmospheres for which is the presence of two photospheres, one for X-mode and the other for O-mode photons. The X-mode photosphere is located deeper in the atmosphere, and therefore, at least if the NSs is modeled as a passive cooler without heat deposition in its external layers, its temperature is higher than that of the O-mode photosphere. As a result, such an atmosphere should radiate mainly in the X-mode, and its emergent radiation should be almost completely linearly polarized.
@@
However, the possibility that the radiating surface of magnetars is described by a classical semi-infinite atmosphere without its own energy sources is still strongly debated [see e.g. 31]. Possibly, energy can be released throughout the entire thickness of the atmosphere as a result of magnetic field dissipation. Heat deposition in the outer layers by returning magnetospheric currents can also largely affect polarization [75, 76]. eXTP has the revolutionary capability to observe the polarization of thermal radiation from a large number of magnetars, extending the preliminary finding by IXPE, and this will unveil the properties of the radiating surface responsible for the thermal component in the spectra of magnetars.
@@
In the presence of a strong magnetic field, the NS atmosphere may turn into a condensed state. This occurs for sufficiently low temperatures, at which atoms, elongated along the direction of , form molecular chains via covalent bonding [77, 78, 79, 80]. Such elongated atoms settle on the star’s surface in a solid/liquid phase, leaving the NS surface exposed (a phenomenon called “magnetic condensation”). There is in fact theoretical and observational evidence that this may occur in some of the colder and more magnetized NSs, such as magnetars (see the next section and references therein) and possibly the XDINSs (the so-called “Magnificent Seven”) [62, 81]. The discrimination between atmospheric and condensed surface emission can be easily performed by eXTP, since low polarization degree is expected for radiation coming from a condensed surface [, dominated by either O- or X-mode photons according to the geometry of view, 28, 31]. Another exciting potentiality of eXTP would be the detection of polarization of X-ray emission from CCOs, since it could resolve the problem of the chemical composition of the atmospheres of some of these non-pulsating objects, such as those in the supernova remnants Cas A [82] and HESS J1731–347 [83]. When these sources have been modelled by making the simple assumption that temperature is uniform on the surface, some authors obtained the conclusion that their atmospheres are made of carbon [84], and, in the case of HESS J1731–347, this led to the estimate of an extremely low mass of the NS, [85]. On the other hand, the assumption of constant temperature is probably unrealistic, in particular in presence of high magnetic fields: the presence of magnetised envelopes can effectively introduce anisotropies in the heat distribution and lead to complex thermal maps [86, 87]. In fact, an alternative and intriguing scenario is that these NSs host instead very high magnetar-like magnetic fields buried by fall-back accretion [see e.g. 88], in which case their surfaces may exhibit hot spots covered by hydrogen atmospheres. In that case, the absence of pulsations may be attributed to a very unfavorable viewing geometry [89]. As discussed in [90], for carbon atmospheres the expected polarization is not detectable, while detecting even a modest degree of polarization (up to a few percent for the CCO in HESS J1731-347, in the 2-8 keV band) would support the hot spot hydrogen atmosphere hypothesis. These authors estimated that measuring this polarization with IXPE required a long exposure time of a few million of seconds, while eXTP, with his enhanced sensitivity, would be excellently placed to make the detection.
@@ 2.2.3 Spectral-continuum polarimetry in magnetars
The nature of the thermal component remains an open problem. Given the ultra-strong magnetic fields of these objects, emission from the star surface may either emerge from a magnetized atmosphere (possibly bombarded by the returning magnetospheric currents) or from a “bare” condensed surface. Despite the fact that the spectrum emitted in the two cases is expected to be not much different (and close to a blackbody distribution), polarization properties are poles apart [31]. As mentioned above, a much lower polarization degree is characteristic of radiation from a condensed surface, compared to the case of a magnetized atmosphere, for which a high polarization degree is generally expected, even when QED mode conversion is accounted for. Because for magnetar-like magnetic fields (), the energy at which mode switching occurs is outside the – band, the working energy range for GPD-based polarimeters like IXPE and eXTP [74].
@@
Different behaviors have been observed in the AXPs 1RXS J170849.0400910 and 1E 1841045. Although the relatively low degree of polarization at low energies (–) still argues in favor of a condensed surface origin, it increases monotonically with energy, up to – at –. Moreover, contrary to the case of 4U 014261, the polarization angle is basically constant throughout the – range. In 1E 1841045, a simultaneous observation with NuSTAR allowed us to associate the high degree of polarization at high energies with a power-law component, which still dominates at [15, 16]. Such a broad-band analysis was not possible for 1RXS J170849.0400910, making it difficult to understand whether the observed polarization properties can only be attributed to surface / magnetospheric emission. Once more, eXTP observations of these sources (with both the SFA and PFA instruments) will provide invaluable insight, helping to fully characterize the spectral continuum of magnetar sources.
@@ 2.3.1 RVM and beyond
Using the RVM, one can obtain important information about the source geometry, mainly the magnetic obliquity (i.e. the angle between the magnetic axis and rotational axis) and the inclination angle between the line of sight and rotational axis, as well as the position angle of the rotational axis projected onto the plane of sky. In each phase, the observer sees different position angles (PAs, position angle of the rotating “vector” relative to the meridian plane defined by the rotational axis and line of sight). Therefore, compared to the case of a NS endowed with a dipolar magnetic field, the RVM will be modified if (i) the “vector” intrinsically changes with time (e.g. due to forced or free precession) or (ii) the field topology deviates from a dipole, e.g. because of the presence of a toroidal component (a twist). A variable PA is often seen in the case of magnetar radio emission[99, 100]. PAs that can not be modeled by the simple form of RVM are also reported in pulsars and FRBs [97, 101]. They all point to a dynamic magnetosphere with a magnetic field geometry more complex than the simple dipole field.
The geometry of a rotating dipole field is shown in Fig. 7. Using simple arguments from spherical trigonometry, it is possible to obtain PA as a function of phase [see, e.g., 36]
@@ 2.3.1.1 Change of the geometry due to precession
In practice, a simpler and straightforward approach can be used by fitting the observed polarization angle with the standard RVM model at different epochs. Then, different sets of RVM parameters will be obtained. In the case of forced precession, the rotation axis of the NS experiences a significant change, due to gravity effects produced by a companion star or a fallback disc. Then, the magnetic axis is dragged together with the spin, so that remains constant. In contrast, free precession arises because of an aspherical deformation of the star, possibly produced by the strong -field. In this case the rotation axis is practically fixed in space (which is approximately coincide with the star’s angular momentum), so that and constant, while the inclination of the magnetic axis varies in time, and a consequence also changes.
@@ 2.3.1.2 Change of field line geometry due to the twist of the magnetic field
If X-ray polarization observations of magnetars are performed during an outburst, a monotonically evolving PA (especially and ) is expected. This is shown in Fig. 9, where the polarization angle resulting from different eXTP simulations is shown [see 27, 104]. Four stages of an outburst have been considered, characterized by different values of the flux – (unabsorbed) and of the radial index of the globally twisted field [see 92]. For simplicity, the source emission is modeled assuming a BB distribution (at a temperature , constant across the surface), polarized in the X-mode. According to the results shown in Fig. 9, the deviation from the classical RVM for different magnetic field twists can be clearly resolved by eXTP observations with sufficiently long exposure.
@@ 2.4.1 Glitches
Starquakes in magnetars might trigger FRBs as revealed by some theoretical and statistical works [151, 152, 150, 153, 154]. The energy functions of FRBs has a universal break which is similar to that of the frequency–magnitude relationship of earthquakes. Such break and the change in power-law indices across this transition can be explained by magnetar starquake models [155]. The energy released during the starquakes is estimated to support X-ray bursts via resonant inverse Compton scattering [156]. The quake model predicts that FRBs will be associated with glitches and quasi-periodic oscillations (QPOs), which are expected to be found in the associated X-ray bursts [157]. Currently, several FRB-like bursts from SGR J1935+2154 are found to be associated with the source spin changes. [158] found a glitch candidate before FRB 200428, the fast rise of this glitch was followed by a delayed spin-up rise, similar to that observed in the Crab pulsar [159]. It is reported that two neighboring glitches within a time interval of approximately nine hours with great details, and the two glitches bracket an FRB on 2022 October 14 [160]. Rapid spin down phase was observed between the two glitches. Previously, [161] reported a spin-down glitch around 2020 October 5 that occurred before three FRB-like bursts, and the observational cadence seems to be much lower than that in [160]. The possibility that this spin-down glitch is actually a glitch followed by an enhanced spin-down rate (just as what was observed by Hu et al. [160]) is worthy of consideration. Note that, apart from starquakes, other models such as tidal capture of an asteroid by a magnetar could also result in the observed spin changes and FRB-like bursts simultaneously [162]. Given the apparent association between glitches/spin-down glitches and galactic FRBs, quick responses and high-cadence observations around the epoch of FRBs should be conducted. If the galactic FRB and glitch/spin-down glitch association is confirmed to be a necessity rather than a coincidence by eXTP, then a comprehensive research on the physical origin of these glitches/spin-down glitches through flux, pulse profile, polarization position angle variations and so on may shed light on the physical origin of galactic FRBs.
@@ 2.4.2.1 Bursting activity and outbursts in magnetar sources
One of the most distinctive behaviors of magnetars is the occurrence of events in which huge amounts of energy are released over different timescales, which, in many cases, led to the discovery of new magnetar sources. Persistent magnetars are characterized by the emission of short (–) bursts, with luminosities in the range –, and intermediate (–) flares, with peak luminosities ranging from to [30]. The rarest giant flares, with luminosities up to in the short (–) initial spike and decreasing down to in the pulsating tail (modulated at the spin frequency of the star), have been definitively observed from only three magnetars [SGR 052666, SGR 190014 and SGR 180620, see 166, 167, 168, 169]. Several magnetar giant flare candidates have been reported from nearby galaxies, such as GRB 200415A in NGC 253 [170]. Conversely, outbursts are typically identified in transient sources following the detection of X-ray bursts by the wide-field instruments on-board such as Swift, Fermi, GECAM missions. These events are characterized by a sudden flux enhancement (by a factor of –), followed by a return to quiescence over timescales ranging from months to several years [171]. In particular, the recent launch of the Einstein Probe and SVOM missions has significantly enhanced our ability to detect new magnetar outbursts [172, 173]. With its large FOV, accurate localization and coverage of the hard X-ray band, eXTP/W2C will be powerful to monitor the burst activity of known magnetars and discover new magnetar sources either in our galaxy or nearby galaxies.
@@
Observing an outburst with eXTP, with its capability of performing simultaneous phase-resolved spectral and polarization measurements, will enable the extraction of extensive information regarding aspects such as the activation mechanism of these phenomena, the energy source driving their long-term emission, and the potential correlations with the emission of short bursts. At the onset of the event, a spectral hardening is usually observed, followed by a slow decay phase in which the X-ray spectral shape softens [179, 29, 180, 181, 182]. In the spectral observations carried out so far, radiation near the peak of an outburst is typically observed to be emitted from large regions, where, due to heat deposition, temperatures are expected to be relatively higher than those of the rest of the surface. This situation represents the most favorable condition to probe the effects of vacuum birefringence (see §2.1). In fact, a sufficiently high surface temperature in the emitting region would prevent magnetic condensation of the gaseous NS atmosphere to occur, so that the intrinsic polarization of emitted radiation can reasonably be expected to be very high. Moreover, if vacuum birefringence were not effective, the observed polarization degree of radiation coming from such an extended and hot cap should be much lower than the intrinsic polarization at the emission, due to the geometrical depolarization caused by tangled magnetic field topology at the surface (see Fig. 3). Hence, measuring a high degree of polarization at the onset of the outburst would provide long-awaited evidence of vacuum birefringence. Furthermore, as the emitting cap shrinks in the decay phase, an increase of the observed polarization degree should be observed if QED effects were not at work, due to geometrical depolarization being less effective for emitting region sizes smaller than (see again Fig. 3). Thanks to the increased effective area and much shorter repointing time compared to those of IXPE, eXTP can map both spectral and polarization properties during a typical magnetar outburst, from the peak phase to the afterglow, so as to put a definitive end to the long-standing problem of the observational test of vacuum birefringence.